# DiffECG: A Versatile Probabilistic Diffusion Model for ECG Signals Synthesis

### v2

@@ 4.3 Experiments and results
Table 1: Classification results for the three baselines for the different settings.
Acharya et al. (2017) Kumar et al. (2019) Kachuee et al. (2018) Accuracy Precision Recall F1 score Accuracy Precision Recall F1 score Accuracy Precision Recall F1 score Setting 1 0.97 0.93 0.89 0.91 0.98 0.87 0.82 0.84 0.96 0.87 0.74 0.77 Setting 2 0.98 0.94 0.91 0.92 0.98 0.93 0.91 0.92 0.97 0.87 0.79 0.82 Setting 3 0.98 0.93 0.93 0.92 0.98 0.93 0.93 0.93 0.98 0.90 0.92 0.91 Setting 4 0.98 0.95 0.93 0.94 0.98 0.96 0.94 0.95 0.99 0.96 0.95 0.95 Setting 5 0.99 0.97 0.95 0.94 0.99 0.97 0.95 0.96 0.99 0.96 0.96 0.96 Setting 6 0.99 0.95 0.93 0.94 0.99 0.96 0.95 0.95 0.99 0.95 0.96 0.96
4.3.1 Quantitative evaluation
Classification evaluation:
In this evaluation, we used three state-of-the-art arrhythmia classification baselines Kumar et al. (2019); Kachuee et al. (2018); Acharya et al. (2017) to assess the impact of adding synthetic data obtained by different generation methods. As a first generation baseline, we adapted the seminal GAN work Goodfellow et al. (2014) to our context. We considered also a state-of-the-art standard GAN approach Delaney et al. (2019) and two other recent GAN-based ECG generation methods Neifar et al. (2022b, a) that successfully incorporate shape prior to guide the generation process leading to prominent results.
The arrhythmia classification baselines Kumar et al. (2019); Kachuee et al. (2018); Acharya et al. (2017) were trained following these five settings:
@@ 4.3.1 Quantitative evaluation
  • Setting 2: + synthetic ECG signals generated by Goodfellow et al. (2014).
  • Setting 3: + synthetic ECG signals generated by Delaney et al. (2019).
  • Setting 4: + synthetic ECG signals generated by Neifar et al. (2022b).
  • Setting 5: + synthetic ECG signals generated by Neifar et al. (2022a).
@@
Tables 1 shows the obtained classification results for the different settings. We can observe that adding synthetic ECG in the training phase systematically improves the arrhythmia classification performances of the three baselines. Moreover, our generation method outperforms the standard GANs Goodfellow et al. (2014); Delaney et al. (2019) and Neifar et al. (2022b) and achieves comparable results to the most advanced GAN-based approach Neifar et al. (2022a). It is worth to notice that these approaches are used only for data generation. However, our method is a generalized approach adapted for generation, imputation and forecasting without any fine-tuning and reconfiguration.
Table 2: Obtained results of quantitative metrics for generation task.
RMSE MAE FID DTW EMD MMD Class N Goodfellow et al. (2014) 1.86e-3 1.47e-3 2.14e-2 29.05 2.35e-2 0.49 Delaney et al. (2019) 1.85e-3 1.42e-3 2.20e-2 28.91 2.31e-2 0.56 Neifar et al. (2022b) 1.83e-3 1.33e-3 2.10e-2 25.67 2.03e-2 0.41 Neifar et al. (2022a) 1.66e-3 1.25e-3 1.74e-2 24.42 1.84e-2 0.40 Ours 1.68e-3 1.20e-3 1.29e-2 22.10 1.54e-2 0.40 Class V Goodfellow et al. (2014) 3.87e-3 2.82e-3 5.03e-2 28.10 4.96e-2 0.60 Delaney et al. (2019) 3.79e-3 2.78e-3 4.98e-2 27.62 4.89e-2 0.59 Neifar et al. (2022b) 3.54e-3 2.64e-3 3.18e-2 24.04 3.52e-2 0.35 Neifar et al. (2022a) 3.02e-3 2.22e-3 1.99e-2 21.20 2.52e-2 0.21 Ours 3.32e-3 2.35e-3 2.61e-02 22.03 2.87e-2 0.25 Class F Goodfellow et al. (2014) 5.37e-3 4.18e-3 8.22e-2 8.63 8.26 0.42 Delaney et al. (2019) 5.49e-3 4.07e-3 6.18e-2 8.35 6.71e-2 0.44 Neifar et al. (2022b) 5.16e-3 3.85e-3 4.48e-2 7.74e-3 5.45e-2 0.40 Neifar et al. (2022a) 4.84e-3 3.47e-3 2.96e-2 6.61 4.23e-2 0.25 Ours 4.97e-03 3.35e-3 3.97e-2 6.50 3.93e-2 0.26
Evaluation metrics:
Following previous studies such as Hazra & Byun (2020); Zhu et al. (2019); Xu et al. (2018), we consider the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), Fréchet Inception Distance (FID), Dynamic Time Warping (DTW), Earth Mover’s Distance (EMD) and Maximum Mean Discrepency (MMD) as performance metrics.
@@
Table 2 reports the values of the used metrics for our ECG generation approach and GAN-based approaches Goodfellow et al. (2014); Delaney et al. (2019); Neifar et al. (2022b, a) on a set of samples from both real data from the test set and synthetic. For all metrics, lower scores mean good results. Overall, our method produces results that are competitive with the four other competing generation methods. Indeed, for all heartbeat classes and metrics, our generation method clearly outperforms the standard GANs generation baseline. In addition, our approach outperforms the advanced GAN Neifar et al. (2022b) for the three classes on all metrics. On the other hand, we achieved better results on some metrics and obtained comparable results on other metrics with the most advanced GAN Neifar et al. (2022a), across the 3 classes. This could be explained by the fact of incorporating advanced prior knowledge about ECG complex dynamic in the generation process in Neifar et al. (2022a). In instance, for class N, we obtained (22.10, 1.54e-2) for (DTW, EMD); while Neifar et al. (2022b) obtained (24.42, 1.84e-2). For example, we obtained for RMSE (1.68e-3, 3.32e-3, 4.97e-3) for classes (N, V, and F) which is comparable to (1.66e-3, 3.02e-3, 4.84e-3) obtained by Neifar et al. (2022a).
@@
Table 3: Obtained results for the imputation task.
RMSE MAE FID EMD MMD Class N LSTM 6.90e-4 5.59e-4 9.88e-3 1.00e-2 0.61 VAE 8.58e-4 5.76e-4 1.40e-2 1.35e-2 0.81 Xu et al. (2022) 6.51e-4 4.62e-4 5.24e-3 5.90e-3 0.32 Ours 3.33e-4 2.215e-4 2.28e-3 2.35e-3 0.24 Class V LSTM 3.62e-03 2.09e-03 4.00e-02 4.47e-2 0.68 VAE 3.51e-3 2.06e-3 5.78e-2 5.64e-2 1.02 Xu et al. (2022) 3.74e-3 2.28e-3 4.04e-2 4.25e-2 0.56 Ours 1.18e-03 7.71e-04 7.13e-03 8.45e-3 0.14 Class F LSTM 4.93e-03 2.80e-3 7.23e-2 6.75e-2 0.53 VAE 5.09e-2 2.54e-3 6.20e-2 5.11e-2 0.28 Xu et al. (2022) 4.86e-3 3.24e-3 3.26e-2 3.45e-2 0.12 Ours 4.09e-03 3.53e-3 3.57e-2 3.46e-2 0.27
@@
Table 4: Obtained results for the forecasting task.
RMSE MAE FID EMD MMD Class N LSTM 8.57e-4 5.52e-4 7.87e-3 6.96 e-3 0.30 VAE 8.87e-4 5.93e-4 1.42e-2 1.38e-2 0.81 Xu et al. (2022) 9.35e-4 6.42e-4 1.47e-2 1.44e-2 0.86 Ours 3.49e-4 2.57e-4 2.22e-3 2.65e-3 0.26 Class V LSTM 4.03e-3 2.44e-3 5.43e-2 5.58e-2 0.87 VAE 4.17e-3 2.36e-3 6.29e-2 6.30e-2 0.99 Xu et al. (2022) 3.63e-3 2.03e-3 3.38e-2 3.75e-2 0.48 Ours 1.07e-03 6.99e-04 6.22e-03 6.92e-3 0.11 Class F LSTM 5.39e-2 4.55e-3 7.33e-2 7.21e-2 0.57 VAE 5.29e-2 3.37e-3 7.64e-2 7.06e-2 0.53 Xu et al. (2022) 4.78e-3 3.18e-3 3.66e-2 3.69e-2 0.14 Ours 3.83e-03 2.77e-3 2.74e-2 3.17e-2 0.27
@@
Figure 2: Examples of synthetic heartbeats for classes N, V, and F classes obtained from our generation approach (a), Neifar et al. (2022a) (b), Neifar et al. (2022b) (c), Goodfellow et al. (2014) (d) and Delaney et al. (2019) (e). The gray background represents the distribution of the real dataset, while the red signals depicts the generated heartbeats.
@@
Figure 3: Examples of synthetic heartbeats for classes N, V, and F classes obtained from our generation approach (a), LSTM (b), VAE (c), and Xu et al. (2022) (d). The gray background represents the distribution of the real dataset, while the red portions depicts the completed portions of the heartbeats.
@@
Figure 4: Examples of synthetic heartbeats for classes N, V, and F classes obtained from our generation approach (a), LSTM (b), VAE (c), and Xu et al. (2022) (d). The gray background represents the distribution of the real dataset, while the red and green portions depicts the completed and ground-truth the heartbeats.
@@
For the evaluation of our approach in the imputation and forecasting tasks, we compared our method to Long short-term memory (LSTM) and Variational auto-encoder (VAE) baselines in addition to a stat-of-the-art method Xu et al. (2022) based on bottleneck dilated convolutional (BDC) self-Attention architecture. Tables 3 and 4 show the obtained results of the used metrics for the imputation and forecasting tasks, respectively. We can observe that our method achieves better results globally for all metrics in both imputation and forecasting tasks. Indeed, our approach outperformed the three selected baselines in the imputation and forecasting tasks across all the metrics in classes N and V. For example, we obtained lower FID in the imputation task (2.28e-3, 7.13e-3) in classes N and V, while ((9.88e-3, 4.00e-2), (1.40e-2,5.78e-2), (5.24e-3, 4.04e-2)) in LSTM, VAE and Xu et al. (2022), respectively. On the other hand, we obtained superior results on RMSE and FID metrics for class F (i.e., minority class) in the imputation task, while obtaining comparable results for the other metrics. Additionally, our method outperformed the other baselines on all metrics in the forecasting task except the MMD metric. For instance, we achieved 0.27 as MMD; whereas 0.14 by Xu et al. (2022). These results demonstrate the effectiveness of our approach in handling missing data and making accurate forecasts for ECG signals.
4.3.2 Qualitative evaluation
@@
For the qualitative assessment, we visually compare the synthesized signals in the three considered scenarios. We select random heartbeats obtained by our approach and the selected baselines and compare them with the distribution of the training dataset. Fig.2 shows examples of synthetic heartbeats from classes (N, V, and F) obtained from our generation approach, Neifar et al. (2022a), Neifar et al. (2022b), Goodfellow et al. (2014) and Delaney et al. (2019) alongside the real distribution of these classes. Generated heartbeats from our approach and the advanced GAN methods exhibit realistic morphology and closely follow the real distribution. The generated heartbeats from standard GANs do not exhibit the complete ECG morphology and are more noisy than other synthetic beats. However, the beats generated by Neifar et al. (2022b) are quite noisier compared to the beats obtained from our method and Neifar et al. (2022a).
@@
Fig. 3 present examples of heartbeats with missing values in the three classes. The completed part of heartbeats by the LSTM (b) VAE (c) and Xu et al. (2022) (d) are overly smoothed compared to completed beats by our approach (a) which is not representative of real ECG signals. Our method accurately completes the missing values in these heartbeats, demonstrating a high imputation performance. Additionally, Fig. 4 displays examples of heartbeats forecasting for classes (N, V, and F) using our approach and the selected baselines. The generated beats from LSTM (b), VAE (c) and Xu et al. (2022) (d) do not always follow the ground-truth heartbeats. For example, the generated heartbeat from class V (c) does not contain the same waves as the ground-truth beat. On the other hand, Fig. 4 (a) shows our approach’s ability to accurately forecast ECG signals with realistic morphology.
5 Conclusion

### v3

@@ IV-C Experiments and results
TABLE I: Classification results for the three baselines for the different settings. Acc., Pr., Re., F1.sc. stand for Accuracy, Precision, Recall, and F1 score.
[22] [23] [24]
Acc. Pr. Re. F1.sc. Acc. Pr. Re. F1.sc. Acc. Pr. Re. F1.sc.
Setting 1 0.97 0.93 0.89 0.91 0.98 0.87 0.82 0.84 0.96 0.87 0.74 0.77
Setting 2 0.98 0.94 0.91 0.92 0.98 0.93 0.91 0.92 0.97 0.87 0.79 0.82
Setting 3 0.98 0.93 0.93 0.92 0.98 0.93 0.93 0.93 0.98 0.90 0.92 0.91
Setting 4 0.98 0.95 0.93 0.94 0.98 0.96 0.94 0.95 0.99 0.96 0.95 0.95
Setting 5 0.99 0.97 0.95 0.94 0.99 0.97 0.95 0.96 0.99 0.96 0.96 0.96
Setting 6 0.99 0.95 0.93 0.94 0.99 0.96 0.95 0.95 0.99 0.95 0.96 0.96
IV-C1 Quantitative evaluation
Classification evaluation
In this evaluation, we used three state-of-the-art arrhythmia classification baselines [23, 24, 22] to assess the impact of adding synthetic data obtained by different generation methods. As a first generation baseline, we adapted the seminal GAN work [9] to our context. We considered also a state-of-the-art standard GAN approach [25] and two other recent GAN-based ECG generation methods [5, 6] that successfully incorporate shape prior to guide the generation process leading to prominent results.
The arrhythmia classification baselines [23, 24, 22] were trained following these five settings:
@@ IV-C1 Quantitative evaluation
  • Setting 2: + synthetic ECG signals generated by [9].
  • Setting 3: + synthetic ECG signals generated by [25].
  • Setting 4: + synthetic ECG signals generated by [5].
  • Setting 5: + synthetic ECG signals generated by [6].
@@
Table I show the obtained classification results for the different settings. We can observe that adding synthetic ECG in the training phase systematically improves the arrhythmia classification performances of the three baselines. Moreover, our generation method outperforms the standard GANs [9, 25] and [5] and achieves comparable results to the most advanced GAN-based approach [6]. It is worth noticing that these approaches are used only for data generation. However, our method is a generalized approach adapted for generation, imputation, and forecasting without any fine-tuning and reconfiguration.
TABLE II: Obtained results of the used metrics across classes (N, V, and F) for generation task.
RMSE MAE FID DTW EMD MMD
N [9] 1.86e-3 1.47e-3 2.14e-2 29.05 2.35e-2 0.49
[25] 1.85e-3 1.42e-3 2.20e-2 28.91 2.31e-2 0.56
[5] 1.83e-3 1.33e-3 2.10e-2 25.67 2.03e-2 0.41
[6] 1.66e-3 1.25e-3 1.74e-2 24.42 1.84e-2 0.40
Ours 1.68e-3 1.20e-3 1.29e-2 22.10 1.54e-2 0.40
V [9] 3.87e-3 2.82e-3 5.03e-2 28.10 4.96e-2 0.60
[25] 3.79e-3 2.78e-3 4.98e-2 27.62 4.89e-2 0.59
[5] 3.54e-3 2.64e-3 3.18e-2 24.04 3.52e-2 0.35
[6] 3.02e-3 2.22e-3 1.99e-2 21.20 2.52e-2 0.21
Ours 3.32e-3 2.35e-3 2.61e-2 22.03 2.87e-2 0.25
F [9] 5.37e-3 4.18e-3 8.22e-2 8.63 8.26e-2 0.42
[25] 5.49e-3 4.07e-3 6.18e-2 8.35 6.71e-2 0.44
[5] 5.16e-3 3.85e-3 4.48e-2 7.74 5.45e-2 0.40
[6] 4.84e-3 3.47e-3 2.96e-2 6.61 4.23e-2 0.25
Ours 4.97e-3 3.35e-3 3.97e-2 6.50 3.93e-2 0.26
Evaluation metrics
Following previous studies such as [26, 27, 28], we consider the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), Fréchet Inception Distance (FID), Dynamic Time Warping (DTW), Earth Mover’s Distance (EMD) and Maximum Mean Discrepancy (MMD) as performance metrics.
@@
Table II reports the values of the used metrics for our ECG generation approach and GAN-based approaches [9, 25, 5, 6] on a set of real ECG from the test set and synthetic ECG. For all metrics, lower scores mean good results. Overall, our method produces results that are competitive with the four other competing generation methods. Indeed, for all heartbeat classes and metrics, our generation method outperforms the standard GANs generation baseline. In addition, our approach outperforms the advanced GAN [5] for the three classes on all metrics. On the other hand, we achieved better results on some metrics and obtained comparable results on other metrics with the most advanced GAN [6], across the 3 classes. This could be explained by the fact of incorporating advanced prior knowledge about ECG complex dynamics in the generation process in [6]. For instance, for class N, we obtained (22.10, 1.54e-2) for (DTW, EMD); while [6] obtained (24.42, 1.84e-2). For example, we obtained for RMSE (1.68e-3, 3.32e-3, 4.97e-3) for classes (N, V, and F) which is comparable to (1.66e-3, 3.02e-3, 4.84e-3) obtained by [6].
@@
TABLE III: Obtained results of the used metrics across classes (N, V, and F) for the imputation task.
RMSE MAE FID EMD MMD
N LSTM 6.90e-4 5.59e-4 9.88e-3 1.00e-2 0.61
VAE 8.58e-4 5.76e-4 1.40e-2 1.35e-2 0.81
[29] 6.51e-4 4.62e-4 5.24e-3 5.90e-3 0.32
Ours 3.33e-4 2.21e-4 2.28e-3 2.35e-3 0.24
V LSTM 3.62e-3 2.09e-3 4.00e-2 4.47e-2 0.68
VAE 3.51e-3 2.06e-3 5.78e-2 5.64e-2 1.02
[29] 3.74e-3 2.28e-3 4.04e-2 4.25e-2 0.56
Ours 1.18e-3 7.71e-4 7.13e-3 8.45e-3 0.14
F LSTM 4.93e-3 2.80e-3 7.23e-2 6.75e-2 0.53
VAE 5.09e-2 2.54e-3 6.20e-2 5.11e-2 0.28
[29] 4.86e-3 3.24e-3 3.26e-2 3.45e-2 0.12
Ours 4.09e-3 3.53e-3 3.57e-2 3.46e-2 0.27
@@
TABLE IV: Obtained results of the used metrics across classes (N, V, and F) for the forecasting task.
RMSE MAE FID EMD MMD
N LSTM 8.57e-4 5.52e-4 7.87e-3 6.96e-3 0.30
VAE 8.87e-4 5.93e-4 1.42e-2 1.38e-2 0.81
[29] 9.35e-4 6.42e-4 1.47e-2 1.44e-2 0.86
Ours 3.49e-4 2.57e-4 2.22e-3 2.65e-3 0.26
V LSTM 4.03e-3 2.44e-3 5.43e-2 5.58e-2 0.87
VAE 4.17e-3 2.36e-3 6.29e-2 6.30e-2 0.99
[29] 3.63e-3 2.03e-3 3.38e-2 3.75e-2 0.48
Ours 1.07e-3 6.99e-4 6.22e-3 6.92e-3 0.11
F LSTM 5.39e-2 4.55e-3 7.33e-2 7.21e-2 0.57
VAE 5.29e-2 3.37e-3 7.64e-2 7.06e-2 0.53
[29] 4.78e-3 3.18e-3 3.66e-2 3.69e-2 0.14
Ours 3.83e-3 2.77e-3 2.74e-2 3.17e-2 0.27
@@
Figure 2: Examples of synthetic heartbeats for classes N, V, and F classes obtained from our generation approach (a), [6] (b), [5] (c), [9] (d) and [25] (e). The gray background represents the distribution of the real dataset, while the red signals depict the generated heartbeats.
@@
Figure 3: Examples of synthetic heartbeats for classes N, V, and F classes obtained from our generation approach (a), LSTM (b), VAE (c), and [29] (d). The gray background represents the distribution of the real dataset, while the red portions depict the completed portions of the heartbeats.
@@
Figure 4: Examples of synthetic heartbeats for classes N, V, and F classes obtained from our generation approach (a), LSTM (b), VAE (c), and [29] (d). The gray background represents the distribution of the real dataset, while the red and green portions depict the completed and ground-truth the heartbeats.
@@
For the evaluation of our approach in the imputation and forecasting tasks, we compared our method to Long short-term memory (LSTM) and Variational auto-encoder (VAE) baselines in addition to a stat-of-the-art method [29] based on bottleneck dilated convolutional (BDC) self-Attention architecture. Tables III and IV show the obtained results of the used metrics for the imputation and forecasting tasks, respectively. We can observe that our method achieves better results globally for all metrics in both imputation and forecasting tasks. Indeed, our approach outperformed the three selected baselines in the imputation and forecasting tasks across all the metrics in classes N and V. For example, we obtained lower FID in the imputation task (2.28e-3, 7.13e-3) in classes N and V, while ((9.88e-3, 4.00e-2), (1.40e-2,5.78e-2), (5.24e-3, 4.04e-2)) in LSTM, VAE and [29], respectively. On the other hand, we obtained superior results on RMSE and FID metrics for class F (i.e., minority class) in the imputation task, while obtaining comparable results for the other metrics. Additionally, our method outperformed the other baselines on all metrics in the forecasting task except the MMD metric. For instance, we achieved 0.27 as MMD; whereas 0.14 by [29]. These results demonstrate the effectiveness of our approach in handling missing data and making accurate forecasts for ECG signals.
IV-C2 Qualitative evaluation
@@
For the qualitative assessment, we start by visually comparing the synthesized signals in the three considered scenarios. We select random heartbeats obtained by our approach and the selected baselines and compare them with the distribution of the training dataset. Fig.2 shows examples of synthetic heartbeats from classes (N, V, and F) obtained from our generation approach, [6], [5], [9] and [25] alongside the real distribution of these classes. Generated heartbeats from our approach and the advanced GAN methods exhibit realistic morphology and closely follow the real distribution. The generated heartbeats from standard GANs do not exhibit complete ECG morphology and are more noisy than other synthetic beats. However, the beats generated by [5] are quite noisier than those obtained from our method and [6].
TABLE V: Recognition rate of both real and fake heartbeats as real obtained by the two cardiologists.
Overall Class N Class V Class F
Cardio. 1 $RR_{real}$ 90% 94.1% 85.7% 83.3%
$RR_{fake}$ 100% 100% 100% 100%
Cardio. 2 $RR_{real}$ 96.6% 100% 100% 83.3%
$RR_{fake}$ 96.6% 100% 100% 83.3%
@@
Fig. 3 presents examples of heartbeats with missing values in the three classes. The completed part of heartbeats by the LSTM (b) VAE (c) and [29] (d) are overly smoothed compared to completed beats by our approach (a) which is not representative of real ECG signals. Our method accurately completes the missing values in these heartbeats, demonstrating a high imputation performance. Additionally, Fig. 4 displays examples of heartbeats forecasting for classes (N, V, and F) using our approach and the selected baselines. The generated beats from LSTM (b), VAE (c), and [29] (d) do not always follow the ground-truth heartbeats. For example, the generated heartbeat from class V (c) does not contain the same waves as the ground-truth beat. On the other hand, Fig. 4 (a) shows our approach’s ability to accurately forecast ECG signals with realistic morphology.
The ECG signals were additionally assessed by cardiologists for a qualitative evaluation of our method. Shuffled sets containing 30 real heartbeats and 30 synthetic heartbeats from the different classes (N, V, and F) were presented to two cardiologists. The main goal is to visually differentiate between synthetic and real ECGs to assess the visual appearance of the generated ECG signals. In addition, this evaluation aimed to determine if these heartbeats could be accurately categorized into their appropriate classes, assessing the signals’ class-specific properties. For this purpose, we computed the recognition rates for both real heartbeats classified as real $RR_{real}$ and fake heartbeats classified as real $RR_{fake}$. These two metrics indicate how the generated signals are realistic and how accurately they belong to their classes. Table V presents the obtained recognition rates by the two cardiologists. The cardiologists achieved high recognition rates for both real and synthetic heartbeats across all types. Notably, the first cardiologist classified all synthetic ECGs from the three classes as real ECGs. These results demonstrate the effectiveness of our proposed approach in generating ECG signals with realistic wave morphology.
V Conclusion